OpenStudy (anonymous):

Matt is asked to simplify x + x2 + x3 + 3x. Here is what he does: x + x2 + x3 + 3x = 4x + x2 + x3 = 4x3 + x3 = 5x3 Did Matt do this problem correctly or incorrectly? If he did it correctly, then what math laws did he follow? If he did it incorrectly, then what math laws did he break?

OpenStudy (anonymous):

just need to know the laws he broke

HanAkoSolo (jamierox4ev3r):

he didn't multiply the like terms. for example, you can't add x and x^2 together. So ultimately, his simplification was incorrect

OpenStudy (anonymous):

thank you!

HanAkoSolo (jamierox4ev3r):

np :)

OpenStudy (anonymous):

can you help with a few more plz??

HanAkoSolo (jamierox4ev3r):

sure

OpenStudy (anonymous):

Matt is asked to factor the following: x2 + 3x + 2, but he is stuck after starting the “area” model. His starting point is below. Finish the factoring for him using our area model and describe in words what you are thinking

HanAkoSolo (jamierox4ev3r):

so to factor this… you would start with this: (x+___)(x+___)

HanAkoSolo (jamierox4ev3r):

then you would fill in what the numbers that simultaneously multiply into 2 and add into three. following me so far?

OpenStudy (anonymous):

yes I think so !

HanAkoSolo (jamierox4ev3r):

right…so what are some factors of 2?

HanAkoSolo (jamierox4ev3r):

@honneybee ?

OpenStudy (anonymous):

sorry internet is slow right now and 1and 2 cuz 2 is a prime

HanAkoSolo (jamierox4ev3r):

that's right :) and because 2+1=3, we know that the final answering to the factoring is (x+2)(x+1) to check your work, you can use the FOIL method

OpenStudy (anonymous):

x^2+3x+2

HanAkoSolo (jamierox4ev3r):

yep…and that's the same as what we started with, so that's the correct thing :)

HanAkoSolo (jamierox4ev3r):

do you understand?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

We can write a square root as a power of 0.5 (or one half). For example, (32)^0.5 = 5.66. Try this with your calculator! a. Use this notation of a fractional power and a property of exponents (from Unit 1) to prove the following: This is challenging, but try it! √ a * b = √a * b √

HanAkoSolo (jamierox4ev3r):

so what exactly are you trying to do here?

OpenStudy (anonymous):

sorry wrong one hold on

OpenStudy (anonymous):

What do you think it means when the weather report says, “There is a 30% chance of rain today?” How do you think they calculate that?